# Two-particle calculations with quantics tensor trains - solving the parquet equations

### v1

@@ VI.1 Hubbard atom
Following the outlined iterative parquet scheme (see Fig. 1), we start from the exact fully irreducible vertices $\Lambda_{r}$, set $\Gamma_{r}=\Lambda_{r},F_{d}=U,F_{m}=-U,F_{s}=2U,F_{t}=\Lambda_{t},G=G_{0}$ and use TCI to efficiently compress the data into QTTs. We then iterate the four BSEs, the parquet equation and the SDE in QTT format by means of the discussed MPO operations, which leads to quick convergence of the results for $\beta=U=1$.
@@
Figure 11(a) shows the maximum absolute normalized error in $\Gamma_{d}$ compared to the exact result after 30 iterations of the iterative parquet cycle with a set tolerance of $10^{-10}$ in the initial TCI. We can now disentangle the two sources of error - the finite size of the discrete frequency grid, corresponding to the error in the respective dense grid calculations indicated by dashed lines, and the QTT approximation. The error due to the QTT approximation for a specified maximum bond dimension can be identified as the difference between the QTT and the dense grid results. Remarkably, the QTT errors quickly converge towards the dense grid results, where larger bond dimensions are needed to reach lower errors for larger grids. E.g. in the case of $R=9$ already with ${D_{\mathrm{max}}}=180$, the error from the QTT approximation de facto vanishes leading to the same result as the dense grid calculation. However, the difference is that in the dense grid calculations objects containing $2^{3R}\simeq 1.34\times 10^{8}$ data points need to be stored, while the QTTs stored for these parameters consist only of $\sim 8.5\times 10^{5}$ elements leading to a compression ratio of $\mathcal{O}(10^{2})$, remarkably, without any loss of accuracy.
@@
If we allow for a small loss of accuracy, even more impressive compression ratios can be achieved, while at the same time reaching very low errors. For instance, it can be observed that already at a bond dimension of 100 maximum normalized errors of $\mathcal{O}(10^{-4})$ can be achieved. In the case of $R=11$ this corresponds to a compression ratio of $\mathcal{O}(10^{4})$ leading only to a tiny fraction of the required memory occupation in comparison to the dense grid calculations. Thus, we observe that the outlined framework can be used to efficiently solve the parquet equations in the compressed QTT format.
@@
Figure 11: (a) Maximum absolute normalized error of the iterative parquet scheme with QTCI for the irreducible vertex $\Gamma_{d}$ compared to the exact values for various grid sizes ($2^{3R}$ grid points) in the case of $U=\beta=1$, with respect to the maximum bond dimension ${D_{\mathrm{max}}}$ after 30 iterations (a) and depending on the iteration in the case of ${D_{\mathrm{max}}}=100$ (b) and ${D_{\mathrm{max}}}=200$ (c). Dashed lines indicate the results from the dense grid calculation without QTCI. In (d), the maximum absolute normalized error of $\Gamma_{d}$ is shown close to the first divergence for $\beta=1.55,U=2.3$ and ${D_{\mathrm{max}}}=100$.
@@
In (b) and (c), we show the maximum normalized errors for various grid sizes with respect to the iteration. At ${D_{\mathrm{max}}}=100$ (b) it can be observed that the dense grid results (dashed lines) are only up to $R=7$ exactly reproduced, while the errors of larger grid calculations level off at $\mathcal{O}(10^{-4})$ in the vicinity of each other. Furthermore, we see that for larger grids slightly larger maximum bond dimensions are required to obtain the same level of error. In comparison, at ${D_{\mathrm{max}}}=200$ (c) it can be observed that that the QTT calculations converge towards the resulting dense grid errors also up to $R=9$.
@@
Next, we show that computations can also be performed for a more challenging case with the parameters $\beta=1.55,U=2.3$, which are chosen in the same way as in Ref. [13]. This case is interesting since $\beta U=3.565$ is very close to the point, where the irreducible vertex in the density channel diverges ($\beta U\approx 3.628$[43, 44, 45]. We show the results of these calculations in Fig. 11(d) for a maximum bond dimension ${D_{\mathrm{max}}}=100$, where the maximum normalized error of $\Gamma_{d}$ is shown with respect to the iteration. Since the parameters are very close to the first divergence line, we use a small mixing parameter $\alpha=0.01$, leading to convergence only after 2600 iterations. In accordance with (b), we observe that in the case of $R\leq 6$ the error plateaus at decreasing levels for increasing $R$, which is due to the increased box size reflecting the results from the dense grid calculation. However, for $R\geq 7$ the situation changes, where the leveled off errors are closer together. This means that the error is now governed by the QTT approximation and not by the finite box size anymore. Still, it can be seen that already a maximum bond dimension of 100 is sufficient to reach absolute normalized errors of $\mathcal{O}(10^{-4})$ for larger values of $R$, similar as in (b). Further improving the results can be achieved by increasing the maximum bond dimension.
@@
These calculations were performed on a single 512 GB node on a cluster without parallelization. For the dense grid calculations performing the same iterative parquet cycle was possible only up to $R=9$ due to the exponentially increasing memory demand. In contrast, using the QTT approach, calculations for $R=11$ were easily carried out on this single node without parallelization. This demonstrates the advantage of QTT, where memory occupation and operations scale logarithmically with increasing resolution, in contrast to the rapid growth in memory and computational costs encountered in standard methods. This allows for efficient computation on large grids, providing a significant advantage over dense grid implementations.
@@ VI.2 Single-impurity Anderson model
After solving the parquet equations for the simplified Hubbard atom case, we extend the QTT framework to the more complex SIAM, where a Hubbard atom-like interacting site is coupled to a bath of non-interacting electrons. Using the parquet approximation in which the fully irreducible vertex is approximated by the bare interaction ($\Lambda_{d}=U,\Lambda_{m}=-U,\Lambda_{s}=2U,\Lambda_{t}=0$), we iteratively evaluate the four BSEs (12), the parquet equation (15) and the SDE (17) with QTTs. Starting from $\Gamma_{r}=\Lambda_{r},F_{r}=\Lambda_{r},G=G_{0}$, we decompose the relevant functions into QTTs using TCI and then make use of the discussed MPO operations in order to iteratively solve the parquet equations. To ensure full convergence, we perform 60 iterations with a linear mixing $\alpha=0.4$. The results presented below were obtained for $\beta=10$, $U=1$, $V=2$, $D=10$ and half-filling, i.e. with $\varepsilon_{d}=-U/2$. For these parameters, the SIAM is in the weakly correlated regime, where the parquet approximation still holds. We compare our results with reference data obtained for the parquet approximation with the state-of-the-art parquet equations implementation on large equidistant frequency grids of Ref. [3] using the single- and multi-boson exchange formulation [12].
Fig. 12(b) shows the irreducible vertex $\Gamma_{d}$ at $\omega=0$ calculated for $R=13$ and a maximum bond dimension ${D_{\mathrm{max}}}=200$, which is in good agreement with reference data in (a). In (c), we show the maximum normalized error of $\Gamma_{d}$ obtained from the QTT calculations in comparison to the reference data depending on the maximum bond dimension for various grid sizes. In agreement with the results for the Hubbard atom, it can be observed that the errors of the QTT calculations converge towards the results of the dense grid calculations, which are indicated by dashed lines. Like in the Hubbard atom case, we exactly reproduce the dense grid results at $R=9$ and ${D_{\mathrm{max}}}=180$ leading to a $\mathcal{O}(10^{2})$ compression ratio de facto without any loss in accuracy due to the QTT approximation. Moreover, since these calculations were performed up to $R=13$, very large compression ratios of $\mathcal{O}(10^{5})$ can be reached, while obtaining a maximum normalized error of $\mathcal{O}(10^{-4})$. In (d), the maximum normalized error with respect to the iteration for ${D_{\mathrm{max}}}=200$ can be observed. We show that for larger grids more iterations are needed to converge due to approaching smaller errors.
@@
Figure 12: Irreducible vertex $\Gamma_{d}$ of the reference data (a) compared to the iterative parquet approximation calculations for $R=13,{D_{\mathrm{max}}}=200$ (b) for $\beta=10,U=1,V=2,D=10,\epsilon=10^{-6}$ and a mixing of $0.4$. The maximum normalized error of $\Gamma_{d}$ with respect to the reference data is shown (c) as a function of the maximum bond dimension ${D_{\mathrm{max}}}$ and (d) dependent on the iteration for ${D_{\mathrm{max}}}=200$, where dashed lines indicate the obtained errors from dense grid calculations.
@@
Finally, let us mention that the performed calculations for $R=13$ would correspond to dense grid calculations with multiple objects of the size of $8\times 2^{3\times 13}\simeq 4.4\times 10^{12}$ bytes. Instead of the necessity of engaging multiple nodes and making use of parallelization for the dense grid calculations, using the QTT framework it was possible to perform these calculations on a single node without parallelization. Applying this approach allowed us to obtain maximum normalized errors of $\mathcal{O}(10^{-4})$, while using only a tiny fraction of the memory required for the corresponding dense grid computations reflecting the computational advantage of the QTT approach.
@@ VII Conclusion and outlook
This work represents a large step forward in solving many-body problems with quantum field theory methods in QTT representation. The chosen example, self-consistent solution of parquet equations, is a challenging one, requiring both efficiency in constructing the QTT representation of two-particle vertices and in evaluation of matrix multiplications and variable shifts within this representation. At the same time, the parquet equations for the simplest model, the Hubbard atom, can be solved analytically, allowing for careful benchmarking and assessment of the performance at each step of the solution separately. In this paper we have numerically shown that the QTT representation of the vertex frequency dependence is suitable for solving the parquet equations and that together with TCI it leads to only logarithmic scaling in the grid size and with fourth power in the maximum bond dimension. For the two examples of Hubbard atom and SIAM we observed that the bond dimension of $\sim$100-200 is enough to obtain the solution with high accuracy. For the case of Hubbard atom we see a saturation (or even decrease) of the bond dimension with increasing the inverse temperature $\beta$. We also expect a saturation or only slow growth of the bond dimension with $\beta$ in more general cases, as conjectured in Ref. 29.
Although for explicit testing we have chosen models, where there are no other degrees of freedom than frequencies (no momentum or orbital dependence), the dissection of the parquet equations solver into operations on QTTs: TCI compression and MPO-MPO contractions, is general and the extension to lattice models is straightforward. All results presented here were obtained on a single core with 512GB memory and the grid sizes in each frequency variable were up to $2^{20}$. This high compressibility of the frequency dependence of vertex functions can be in the future exploited to (i) solve parquet equations for lattice systems with high momentum resolution and orbital degrees of freedom, needed to address material properties; (ii) apply QTCI to other vertex based methods, such as ladder extensions of dynamical mean-field theory [37], multiscale single-boson exchange methods [46], functional renormalization group [47], or the Migdal-Eliashberg theory in ab initio calculations [48, 49].
@@ Acknowledgements.
This work was funded in part by the Austrian Science Fund (FWF) Projects No. P 36332 and V 1018 and by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC-2111 (Project No. 390814868). This work is part of the Munich Quantum Valley, supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. H.S. was supported by JSPS KAKENHI Grants No. 21H01041, No. 21H01003, and No. 23H03817 as well as JST FOREST Grant No. JPMJFR2232 and JST PRESTO Grant No. JPMJPR2012, Japan. Calculations have been partly performed on the Vienna Scientific Cluster (VSC).
@@ Appendix A Parquet equation and frequency conventions
The parquet equation (15) gives the full vertex $F$ as a simple sum of the fully irreducible vertex $\Lambda$ and vertices reducible in the ph, pp, and $\overline{ph}$ channels. The reducible vertices however are represented in their ”channel native” frequency parametrization. After applying the parametrization changes and collecting the spin components, the final expressions are linear combinations of different spin components and frequency shifted arguments. For $F_{d}$ see Eq. (16) and for the remaining three spin combinations we have
$F_{m}^{\nu\nu^{\prime}\omega}$ $=\Lambda_{m}^{\nu\nu^{\prime}\omega}+\Phi_{m}^{\nu\nu^{\prime}\omega}-\frac{1}{2}\Phi_{d}^{\nu(\nu+\omega)(\nu^{\prime}-\nu)}+\frac{1}{2}\Phi_{m}^{\nu(\nu+\omega)(\nu^{\prime}-\nu)}$
$-\frac{1}{2}\Phi_{s}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}+\frac{1}{2}\Phi_{t}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}$ (27a)
$F_{s}^{\nu\nu^{\prime}\omega}$ $=\Lambda_{s}^{\nu\nu^{\prime}\omega}+\Phi_{s}^{\nu\nu^{\prime}\omega}+\frac{1}{2}\Phi_{d}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}-\frac{3}{2}\Phi_{m}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}$
$+\frac{1}{2}\Phi_{d}^{\nu(-\nu^{\prime}-\omega)(\nu^{\prime}-\nu)}-\frac{3}{2}\Phi_{m}^{\nu(-\nu^{\prime}-\omega)(\nu^{\prime}-\nu)}$ (27b)
$F_{t}^{\nu\nu^{\prime}\omega}$ $=\Lambda_{t}^{\nu\nu^{\prime}\omega}+\Phi_{t}^{\nu\nu^{\prime}\omega}+\frac{1}{2}\Phi_{d}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}+\frac{1}{2}\Phi_{m}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}$
$-\frac{1}{2}\Phi_{d}^{\nu(-\nu^{\prime}-\omega)(\nu^{\prime}-\nu)}-\frac{1}{2}\Phi_{m}^{\nu(-\nu^{\prime}-\omega)(\nu^{\prime}-\nu)}.$ (27c)
@@
Additionally, we also need a transformation between ph and $\overline{ph}$ representations to obtain $\Phi_{\overline{ph}}$. This transformation exploits the so called crossing symmetry relation between the ph and $\overline{ph}$ frequency channels.

### v2

@@ VI.1 Hubbard atom
Following the iterative parquet scheme outlined in Fig. 1, we start from the exact fully irreducible vertices $\Lambda_{r}$, set $\Gamma_{r}=\Lambda_{r},F_{d}=U,F_{m}=-U,F_{s}=2U,F_{t}=\Lambda_{t},G=G_{0}$ and use TCI to efficiently compress the data into QTTs. We then iterate the four BSEs, the parquet equation and the SDE in QTT format by means of the discussed MPO operations, which leads to quick convergence of the results for $\beta=U=1$.
@@
Figure 11(a) shows the maximum absolute normalized error in $\Gamma_{d}$ compared to the exact result after 30 iterations of the iterative parquet cycle with a set tolerance of $10^{-10}$ in the initial TCI. We can now disentangle the two sources of error—the finite size of the discrete frequency grid, corresponding to the error in the respective dense grid calculations indicated by dashed lines, and the QTT approximation. The error due to the QTT approximation for a specified maximum bond dimension can be identified as the difference between the QTT and the dense grid results. Remarkably, the QTT errors quickly converge towards the dense grid results, where larger bond dimensions are needed to reach lower errors for larger grids. E.g. in the case of $R=9$ already with $D_{\mathrm{max}}=180$, the error from the QTT approximation de facto vanishes leading to the same result as the dense grid calculation. However, the difference is that in the dense grid calculations objects containing $2^{3R}\simeq 1.34\times 10^{8}$ data points need to be stored, while the QTTs stored for these parameters consist only of $\sim 8.5\times 10^{5}$ elements leading to a compression ratio of $\mathcal{O}(10^{2})$, remarkably, without any loss of accuracy.
@@
If we allow for a small loss of accuracy, even more impressive compression ratios can be achieved, while at the same time reaching very low errors. For instance, it can be observed that already at a bond dimension of 100 maximum normalized errors $<10^{-3}$ can be achieved. In the case of $R=11$ this corresponds to a compression ratio of $\mathcal{O}(10^{4})$ leading only to a tiny fraction of the required memory occupation in comparison to the dense grid calculations. Thus, we observe that the outlined framework can be used to efficiently solve the parquet equations in the compressed QTT format.
@@
Figure 11: (a) Maximum absolute normalized error $||\Delta\Gamma_{d}||_{\infty}:=||\Gamma_{d,\text{iterative-parquet}}-\Gamma_{d,\text{exact}}||_{\infty}/||\Gamma_{d,\text{exact}}||_{\infty}$ of the iterative parquet scheme with QTCI for the irreducible vertex $\Gamma_{d}$ compared to the exact values for various grid sizes ($2^{3R}$ grid points) in the case of $U=\beta=1$, plotted (a) as a function of $D_{\mathrm{max}}$ after 30 iterations, and as a function of iteration number for (b) $D_{\mathrm{max}}=100$ and (c) $D_{\mathrm{max}}=200$. Dashed lines indicate the results from the dense grid calculation without QTCI. (d) The maximum absolute normalized error of $\Gamma_{d}$, shown close to the first divergence, for $\beta=1.55,U=2.3$ and $D_{\mathrm{max}}=100$.
@@
In Figs. 11(b) and (c), we show the maximum normalized errors for various grid sizes with respect to the iteration. At $D_{\mathrm{max}}=100$ (b) the dense grid results (dashed lines) are only up to $R=7$ exactly reproduced, while the errors of larger grid calculations level off below $10^{-3}$ in the vicinity of each other. Furthermore, we see that for larger grids slightly larger maximum bond dimensions are required to obtain the same level of error. In comparison, at $D_{\mathrm{max}}=200$ (c) the QTT calculations converge towards the resulting dense grid errors also up to $R=9$.
@@
Next, we show that computations can also be performed for a more challenging case with the parameters $\beta=1.55,U=2.3$, which are chosen in the same way as in Ref. [13]. This case is interesting since $\beta U=3.565$ is very close to the point, where the irreducible vertex in the density channel diverges ($\beta U\approx 3.628$[48, 49, 50, 51, 52]. We show the results of these calculations in Fig. 11(d) for a maximum bond dimension $D_{\mathrm{max}}=100$, where the maximum normalized error of $\Gamma_{d}$ is shown with respect to the iteration. Since the parameters are very close to the first divergence line, we use a small mixing parameter $\alpha=0.01$, leading to convergence only after 2600 iterations. In accordance with (b), we observe that in the case of $R\leq 6$ the error plateaus at decreasing levels for increasing $R$, which is due to the increased box size reflecting the results from the dense grid calculation. However, for $R\geq 7$ the situation changes, where the leveled off errors are closer together. This means that the error is now governed by the QTT approximation and not by the finite box size anymore. Still, it can be seen that already a maximum bond dimension of 100 is sufficient to reach absolute normalized errors $<10^{-3}$ for larger values of $R$, similar as in (b). Further improving the results can be achieved by increasing the maximum bond dimension.
@@
These calculations were performed on a single 512 GB node on a cluster without parallelization, where in Fig. 12, the runtime of a single iteration of the iterative parquet scheme is shown. For the dense grid calculations performing the same iterative parquet cycle was possible only up to $R=9$ due to the exponentially increasing memory demand. In contrast, using the QTT approach, calculations for $R=11$ were easily carried out on this single node without parallelization. This demonstrates the advantage of QTTs, where memory occupation and operations scale logarithmically with increasing resolution (see Fig. 12(b)), in contrast to the rapid growth in memory and computational costs encountered in standard methods. This allows for efficient computation on large grids, providing a significant advantage over dense grid implementations.
Figure 12: Runtime of a single iteration of the iterative parquet scheme for various maximum bond dimensions and grid size parameters $R$ in the case of $\beta=U=1$. The dashed lines indicate (b) the quartic runtime scaling with $D_{\mathrm{max}}$ (governed by the BSEs), and (c) linear increase with $R$.
@@ VI.2 Single-impurity Anderson model
After solving the parquet equations for the simplified Hubbard atom case, we extend the QTT framework to the more complex SIAM, where a Hubbard atom-like interacting site is coupled to a bath of non-interacting electrons. Using the parquet approximation in which the fully irreducible vertex is approximated by the bare interaction ($\Lambda_{d}=U,\Lambda_{m}=-U,\Lambda_{s}=2U,\Lambda_{t}=0$), we iteratively evaluate the four BSEs (12), the parquet equation (15) and the SDE (17) with QTTs. Starting from $\Gamma_{r}=\Lambda_{r},F_{r}=\Lambda_{r},G=G_{0}$, we decompose the relevant functions into QTTs using TCI and then make use of the discussed MPO operations in order to iteratively solve the parquet equations. To ensure full convergence, we perform 60 iterations with a linear mixing $\alpha=0.4$. The results presented below were obtained for $\beta=10$, $U=1$, $V=2$, $D=10$ and half-filling, i.e. with $\varepsilon_{d}=-U/2$. For these parameters, the SIAM is in the weakly correlated regime, where the parquet approximation still holds. We compare our results with reference data obtained for the parquet approximation with the state-of-the-art parquet equations implementation on large equidistant frequency grids of Ref. 3 using the single- and multi-boson exchange formulation [12].
@@
Figure 13: Irreducible vertex $\Gamma_{d}$ of the reference data (a) compared to the iterative parquet approximation calculations for $R=13,D_{\mathrm{max}}=200$ (b) for $\beta=10,U=1,V=2,D=10,\epsilon=10^{-6}$ and a mixing of $0.4$. (c)–(d) The maximum normalized error of $\Gamma_{d}$ ($||\Delta\Gamma_{d}||_{\infty}:=||\Gamma_{d,\mathrm{iterative-parquet}}-\Gamma_{d,\mathrm{ref}}||_{\infty}/||\Gamma_{d,\mathrm{ref}}||_{\infty}$) with respect to the reference data is shown (c) as a function of the maximum bond dimension $D_{\mathrm{max}}$ and (d) dependent on the iteration for $D_{\mathrm{max}}=200$, where dashed lines indicate the obtained errors from dense grid calculations.
Fig. 13(b) shows the irreducible vertex $\Gamma_{d}$ at $\omega=0$ calculated for $R=13$ and a maximum bond dimension $D_{\mathrm{max}}=200$, which is in good agreement with reference data in (a). In (c), we show the maximum normalized error of $\Gamma_{d}$ obtained from the QTT calculations in comparison to the reference data depending on the maximum bond dimension for various grid sizes. In agreement with the results for the Hubbard atom, it can be observed that the errors of the QTT calculations converge towards the results of the dense grid calculations, which are indicated by dashed lines. Like in the Hubbard atom case, we exactly reproduce the dense grid results at $R=9$ and $D_{\mathrm{max}}=180$ leading to a $\mathcal{O}(10^{2})$ compression ratio de facto without any loss in accuracy due to the QTT approximation. Moreover, since these calculations were performed up to $R=13$, very large compression ratios of $\mathcal{O}(10^{5})$ can be reached, while obtaining a maximum normalized error $<10^{-3}$. In (d), the maximum normalized error with respect to the iteration for $D_{\mathrm{max}}=200$ can be observed. We show that for larger grids more iterations are needed to converge due to approaching smaller errors.
@@
Finally, let us mention that the performed calculations for $R=13$ would correspond to dense grid calculations with multiple objects of the size of $8\times 2^{3\times 13}\simeq 4.4\times 10^{12}$ bytes. Instead of the necessity of engaging multiple nodes and making use of parallelization for the dense grid calculations, using the QTT framework it was possible to perform these calculations on a single node without parallelization. Applying this approach allowed us to obtain maximum normalized errors $<10^{-3}$, while using only a tiny fraction of the memory required for the corresponding dense grid computations. This shows the computational advantage of the QTT approach.
@@ VII Conclusion and outlook
This work represents a large step forward in solving many-body problems with quantum field theory methods in QTT representations. The chosen example, the self-consistent solution of parquet equations, is a challenging one, requiring both efficiency in constructing the QTT representation of two-particle vertices and in evaluation of matrix multiplications and variable shifts within this representation. At the same time, the parquet equations for the simplest model, the Hubbard atom, can be solved analytically, allowing for careful benchmarking and assessment of the performance at each step of the solution separately. In this paper we have numerically shown that the QTT representation of the vertex frequency dependence is suitable for solving the parquet equations and that together with TCI it leads to only logarithmic scaling in the grid size and with fourth power in the maximum bond dimension. For the two examples of Hubbard atom and SIAM we observed that the bond dimension of $\sim$100-200 is enough to obtain the solution with high accuracy. For the case of Hubbard atom we see a saturation (or even decrease) of the bond dimension with increasing the inverse temperature $\beta$. We also expect a saturation or only slow growth of the bond dimension with $\beta$ in more general cases, as conjectured in Ref. 31.
Although for explicit testing we have chosen models having no other degrees of freedom than frequencies (no momentum or orbital dependence), the dissection of the parquet equations solver into operations on QTTs—TCI compression and MPO-MPO contractions—is general and the extension to lattice models is straightforward. All results presented here were obtained on a single core with 512 GB memory and the grid sizes in each frequency variable were up to $2^{20}$. This high compressibility of the frequency dependence of vertex functions can in the future be exploited (i) to solve parquet equations for lattice systems with high momentum resolution and orbital degrees of freedom, needed to address material properties; (ii) to apply QTCI to other vertex based methods, such as the functional renormalization group (fRG) [53, 54, 55]; ladder extensions [39] or fRG extensions [56] of dynamical mean-field theory (possibly using as input results for the local vertex obtained using the numerical renormalization group [57, 2, 58]); embedded multi-boson exchange methods [59]; or the Migdal-Eliashberg theory in ab initio calculations [60, 61].
@@ Acknowledgements.
This work was funded in part by the Austrian Science Fund (FWF) Projects No. P 36332 (Grant DOI 10.55776/P36332) and V 1018 (Grant DOI 10.55776/V1018). For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. This work was funded in part by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC-2111 (Project No. 390814868). M.K.R. and J.v.D. acknowledge support from DFG grant LE 3883/2-2. This work is part of the Munich Quantum Valley, supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. H.S. was supported by JSPS KAKENHI Grants No. 21H01041, No. 21H01003, and No. 23H03817 as well as JST FOREST Grant No. JPMJFR2232 and JST PRESTO Grant No. JPMJPR2012, Japan. Calculations have been partly performed on the Vienna Scientific Cluster (VSC).
@@ Appendix A Parquet equation and frequency conventions
The parquet equation (15) gives the full vertex $F$ as a simple sum of the fully irreducible vertex $\Lambda$ and vertices reducible in the ph, pp, and $\overline{\textit{ph}}$ channels. The reducible vertices however are represented in their ”channel native” frequency parametrization. After applying the parametrization changes and collecting the spin components, the final expressions are linear combinations of different spin components and frequency shifted arguments. For $F_{d}$ see Eq. (16) and for the remaining three spin combinations we have
$F_{m}^{\nu\nu^{\prime}\omega}$ $=\Lambda_{m}^{\nu\nu^{\prime}\omega}+\Phi_{m}^{\nu\nu^{\prime}\omega}-\tfrac{1}{2}\Phi_{d}^{\nu(\nu+\omega)(\nu^{\prime}-\nu)}+\tfrac{1}{2}\Phi_{m}^{\nu(\nu+\omega)(\nu^{\prime}-\nu)}-\tfrac{1}{2}\Phi_{s}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}+\tfrac{1}{2}\Phi_{t}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}\,,$ (25a)
$F_{s}^{\nu\nu^{\prime}\omega}$ $=\Lambda_{s}^{\nu\nu^{\prime}\omega}+\Phi_{s}^{\nu\nu^{\prime}\omega}+\tfrac{1}{2}\Phi_{d}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}-\tfrac{3}{2}\Phi_{m}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}+\tfrac{1}{2}\Phi_{d}^{\nu(-\nu^{\prime}-\omega)(\nu^{\prime}-\nu)}-\tfrac{3}{2}\Phi_{m}^{\nu(-\nu^{\prime}-\omega)(\nu^{\prime}-\nu)}\,,$ (25b)
$F_{t}^{\nu\nu^{\prime}\omega}$ $=\Lambda_{t}^{\nu\nu^{\prime}\omega}+\Phi_{t}^{\nu\nu^{\prime}\omega}+\tfrac{1}{2}\Phi_{d}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}+\tfrac{1}{2}\Phi_{m}^{\nu\nu^{\prime}(-\omega-\nu-\nu^{\prime})}-\tfrac{1}{2}\Phi_{d}^{\nu(-\nu^{\prime}-\omega)(\nu^{\prime}-\nu)}-\tfrac{1}{2}\Phi_{m}^{\nu(-\nu^{\prime}-\omega)(\nu^{\prime}-\nu)}\,.$ (25c)
@@
Additionally, we also need a transformation between ph and $\overline{\textit{ph}}$ representations to obtain $\Phi_{\overline{\textit{ph}}}$. This transformation exploits the so called crossing symmetry relation between the ph and $\overline{\textit{ph}}$ frequency channels.